the vapor pressures of pure propyl alcohol and isopropyl alcohol are 21.0 mmHg and 45.2 mmHg. respectively, above a solution in which the mole fraction of isopropyl alcohol is 0.750

1
Expert's answer
2021-01-13T13:55:36-0500

Pressure is directly proportional to Mole

Therefore,

P = kn

k= P/n

P1/n1 = P2/n2

21.0 mmHg / x = 45.2 mmHg/0.750

x = (21.0 mmHg * 0.750) / 45.2 mmHg

x= 0.35 (pure propyl alcohol  mole ratio)